In garden peas a pure line that is tall with purple flowers is crossed to another pure line that is short with white flowers. The F1s are all short and white. You cross the F1s with each other and get 900 short white, 315 short purple, 307 tall white, and 99 tall purple. You allow 90 of the short purple F2 plants to self-fertilize and then select a single seed from each one. What phenotypes would you expect to see in the plants reared from these seeds and in what numbers? Why?
1
Expert's answer
2019-06-05T08:18:57-0400

Let's assume:

A - short

a - tall

B - white

b - purple


Parents:

AABB x aabb

Gamets:

Ab ab

F1:

AaBb - 100% short and white


P2:

AaBb x AaBb

F2:

AABB - 1/16 - short, white

AaBB - 1/8 - short, white

AaBb - 5/16 - short, white

AAbb - 1/16 - short, purple

AABb - 1/16 - short, white

aaBb - 1/8 - tall, white

Aabb - 1/8 - short, purple

aaBB - 1/16 - tall, white

aabb - 1/16 - tall, purple


Total progeny = 900+315+307+99 = 1621


Short white - theoretical 56.25%, practical 55.5%

Short purple - theoretical 18.75%, practical 19.4%

Tall white - theoretical 18.75%, practical 18.9%

Tall purple - theoretical 6.25%, practical 6.1%


So, the assumption was correct.


Since purple is a recessive trait, F3 will be purple in 100%. So, let's analyse the ratio of short and tall plants in F3.

Set of 90 plants consist of 2*Aa and 1*AA

P3.1:

Aa x Aa

G3.1:

A a A a

F3.1:

AA, 2*Aa, aa = 60


P3.2:

AA x AA

G 3.2:

A A

F3.2:

AA = 30


F3 (total):

AA - 45 - short

Aa - 30 - short

aa - 15 - tall


Answer: expected phenotypes:

short purple - 75

tall purple - 15